# Inductive Gradient Adjustment for Spectral bias in implicit neural representations

### v1

@@ Abstract
Implicit Neural Representations (INRs), as a versatile representation paradigm, have achieved success in various computer vision tasks. Due to the spectral bias of the vanilla multi-layer perceptrons (MLPs), existing methods focus on designing MLPs with sophisticated architectures or repurposing training techniques for highly accurate INRs. In this paper, we delve into the linear dynamics model of MLPs and theoretically identify the empirical Neural Tangent Kernel (eNTK) matrix as a reliable link between spectral bias and training dynamics. Based on eNTK matrix, we propose a practical inductive gradient adjustment method, which could purposefully improve the spectral bias via inductive generalization of eNTK-based gradient transformation matrix. We evaluate our method on different INRs tasks with various INR architectures and compare to existing training techniques. The superior representation performance clearly validates the advantage of our proposed method. Armed with our gradient adjustment method, better INRs with more enhanced texture details and sharpened edges can be learned from data by tailored improvements on spectral bias.
@@ 1 Introduction
The main idea of implicit neural representations (INRs) is using neural networks such as multi-layer perceptrons (MLPs) to parameterize discrete signals in an implicit and continuous manner. Benefiting from the continuity and implicit nature, INRs have gained great attention as a versatile signal representation paradigm and achieved state-of-the-art performance across a wide range of computer vision tasks such as signal representation (Sitzmann et al., 2020; Liang et al., 2022; Saragadam et al., 2023; Guo et al., 2023), 3D shape reconstruction (Cai et al., 2024; Shi et al., 2024a; Zhu et al., 2024) and novel view synthesis (Mildenhall et al., 2021; Müller et al., 2022; Fathony et al., 2020).
@@
However, obtaining high-precision INRs is non-trivial, MLPs with ReLU activation function often fail to represent high-frequency details. Such tendency of MLPs to represent simple patterns of target function is referred to as spectral bias (Rahaman et al., 2019; Xu, 2018). To improve the performance of INRs, great efforts have been made to alleviate or overcome this bias of MLPs. These efforts have primarily focused on how MLPs are constructed, such as complex input embeddings (Tancik et al., 2020; Xie et al., 2023; Müller et al., 2022) and sophisticated activation functions (Sitzmann et al., 2020; Saragadam et al., 2023). Recently, two studies find that training modifications based on traditional training techniques, i.e., Fourier reparameterized training (FR) (Shi et al., 2024a) and batch normalization (BN) (Cai et al., 2024), can improve training dynamics, thereby overcoming the spectral bias without altering inference structures.
@@
Despite these works serendipitously discover that such FR and BN training could alleviate the spectral bias and lead to better representation performance, the mechanism behind their improvement is still unclear. Moreover, there is no clear guidance on the choice of Fourier bases matrix for reparameterization or batch normalization layer to confront spectral bias with varying degrees in a variety of INR tasks. Lack of reliable insight and guidance results in suboptimal improvements in their broader applications. Therefore, how to adjust training dynamics of MLPs to purposefully overcome spectral bias remains an open and valuable research challenge.
@@
In this paper, we delve into the linear dynamics model of MLPs and propose an effective training dynamics adjustment strategy guided by theoretical derivation to purposefully alleviate the spectral bias issue. Our strategy allows for tailored impacts on spectral bias based on the spectrum of target signals, thereby achieving more precise INRs. Specifically, these impacts on spectral bias are tailored by purposefully adjusting the spectrum of the Neural Tangent Kernel (NTK) matrix (Jacot et al., 2018), which closely connects the spectral bias with training dynamics. Given that NTK matrix is not available in most cases, we theoretically identify the empirical NTK (eNTK) matrix as a tractable surrogate. Empirical results corroborate this theoretical results. Based on eNTK matrix, we further propose an inductive gradient adjustment (IGA) method which could purposefully improve the spectral bias via inductive generalization of eNTK-based gradient transformation matrix with millions of data points. We give both theoretical and empirical analysis of our method, demonstrating how it tailors impacts on spectral bias. Besides, our IGA method can work with previous structural improvement methods such as positional encoding (Tancik et al., 2020) and periodic activation function (Sitzmann et al., 2020). We validate our method in various vision applications of INRs and compare it to previous training dynamics adjustment methods. Experimental results show the superiority of training dynamics adjustment. Our contributions are summarized as follows:
@@
  • We connect spectral bias with linear dynamics model of MLPs and derive a training dynamics adjustment strategy guided by a theoretical standpoint to purposefully improve the spectral bias, i.e., tailored impacts on spectral bias based on the spectrum of target signals.
@@
  • We provide detailed experimental analyses across a wide range of implicit neural representation tasks. Comparing to previous training dynamics methods, our method allows for tailored improvement on spectral bias of commonly MLPs and provides implicit neural representations with more high-frequency details.
@@
Implicit neural representations. Recently, implicit neural representations (INRs) representing the discrete signal as an implicit continuous function by MLPs have gained lots of attention. Comparing to traditional grid-based signal representation method, INRs have shown the remarkable representation accuracy and memory-efficient property in a wide range of representation tasks such as 1D audio representation (Sitzmann et al., 2020; Kim et al., 2022), 2D image representation (Klocek et al., 2019; Strümpler et al., 2022) , 3D shape representation (Park et al., 2019; Martel et al., 2021), novel view synthesis (Mildenhall et al., 2021; Saragadam et al., 2023) and virtual reality (Deng et al., 2022). However, vanilla MLPs with ReLU activation function fails to represent complex signals. Therefore, various modifications have been studied. One category of modifications focuses on the embeddings of inputs. Mildenhall et al. (2021) find that using positional encoding (PE) as inputs enhances the representation of high-frequency details in neural radiance fields. Further, Takikawa et al. (2021); Martel et al. (2021); Xie et al. (2023) adopt learned features to encode inputs, achieving better INRs. A different category of modifications concentrates on activation functions. Sitzmann et al. (2020) find that periodic functions, such as Sine function, can obtain more accurate representation than ReLU activation function. Sine functions are also explored by Fathony et al. (2020) in multiplicative filter networks. Saragadam et al. (2023) adopt a complex Gabor wavelet activation and achieve robust and accurate representations. Considering a broader framework, the aforementioned work can be summarized as efforts to construct implicit neural representation models or MLPs with more powerful representational capacity (Yüce et al., 2022). Recently, another category, distinct from those previously discussed, focuses on the training process of MLPs. Shi et al. (2024a) find that learning parameters in the Fourier domain, i.e., Fourier reparameterized training (FR), can improve approximation accuracy of INRs without altering the inference structure. Cai et al. (2024) repurpose that the classic batch normalization layer (BN) can also improve performance of INRs.
@@
Spectral bias. Spectral bias typically refers to a learning bias of MLPs that MLPs tend to learn simple patterns of training data or the low-frequency components of the target function. Great efforts have been made to dissect this bias. Rahaman et al. (2019) find that spectral bias varies with the width and depth of MLPs and the complexity of input manifold by experimental and theoretical analyses. Generally, wider or deeper MLPs can learn high frequency information more efficiently, and increasing the complexity of input manifolds serves a similar purpose. Xu (2018) attribute this bias of MLPs with Tanh activation function to the uneven distribution of gradients in the frequency domain through Fourier decomposition. From the perspective of linear dynamics model of MLPs, spectral bias is induced by the uneven distributed eigenvalues of the corresponding Neural Tangent Kernel (NTK) matrix (Jacot et al., 2018; Arora et al., 2019; Ronen et al., 2019). Inspired by this theoretical result, Tancik et al. (2020); Shi et al. (2024a); Cai et al. (2024) try to elucidate spectral bias of different models by observing eigenvalue distribution and find that PE, FR and BN help to attain a more uniform eigenvalue distribution of the NTK matrix. Similarly, Geifman et al. (2023) theoretically discuss impacts of modifying the NTK matrix spectrum in infinitely wide networks.
In this paper, we adopt the insight of attaining more uniform eigenvalue distribution of the NTK matrix to overcome the spectral bias. For the first time, we show that modifying the eNTK matrix spectrum can have the similar impacts on spectral bias and propose a practical inductive gradient adjustment (IGA) method, which improves spectral bias of INRs via inductive generalization of eNTK-based gradient transformation matrix.
@@ 3 Method
In this section, we firstly review the linear dynamics model of MLPs, indicating the potential of adjusting the spectrum of NTK matrix to purposefully improve the spectral bias. Then we analyze the NTK-based adjustment method and show that the intractability of the NTK matrix renders it impractical. By further theoretical analysis, we identify empirical NTK matrix as a tractable proxy, but it succumbs to dimensionality challenges as the data size increases. Therefore, we propose a practical gradient adjustment method, which could purposefully improve the spectral bias via inductive generalization of eNTK-based gradient transformation matrix.
3.1 Background: Connecting spectral bias with training dynamics
Considering a discrete signal $S=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ over $\mathbb{R}^{d_{0}}\times\mathbb{R}^{1}$, the INR of this signal is denoted as $f({\bm{x}},\Theta)$. An MLP with constant width $m$ is adopted to parameterize $f({\bm{x}},\Theta)$. Parameters $\Theta$ are optimized via gradient descent under the supervision of the squared error loss $\mathcal{L}(\mathbf{x};\Theta)=\frac{1}{2}\sum_{i=1}^{N}(f(\mathbf{x}_{i};\Theta)-y_{i})^{2}$. We denote the residual $\left(f({\bm{x}}_{i};\Theta_{t})-y_{i}\right)_{i=1}^{n}$ at time step $t$ as ${\bm{r}}_{t}$. Training dynamics of ${\bm{r}}_{t}$ can be approximated by the linear dynamics model when the width $m$ is large enough and learning rate $\eta$ is small enough (Du et al., 2018; Arora et al., 2019; Lee et al., 2019; Geifman et al., 2023): ${\bm{r}}_{t}=({\bm{I}}-\eta{\bm{K}}){\bm{r}}_{t-1}$, where ${\bm{K}}$ is the NTK matrix on $S$ defined by ${\bm{K}}=\mathbb{E}_{\Theta_{0}}[\nabla_{\Theta}f({\bm{X}};\Theta_{0})^{\top}\nabla_{\Theta}f({\bm{X}};\Theta_{0})]$. With the eigenvalue decomposition of ${\bm{K}}=\sum_{i=1}^{N}\lambda_{i}{\bm{v}}_{i}{\bm{v}}_{i}^{\top}$, training dynamics of ${\bm{r}}_{t}$ can be characterized as follows (Arora et al., 2019):
@@ 3.1 Background: Connecting spectral bias with training dynamics
Equation 1 shows that convergence rate of $f({\bm{x}},\Theta)$ at the projection direction ${\bm{v}}_{i}^{\top}$ with larger $\lambda_{i}$ will be faster. For vanilla MLPs, projection directions related to high frequencies are consistently assigned to small eigenvalues, while those related to low frequencies correspond to larger eigenvalues (Ronen et al., 2019; Bietti & Mairal, 2019; Heckel & Soltanolkotabi, 2020). The uneven spectrum of ${\bm{K}}$ leads to extremely slow convergence to the high frequency components of signal, thereby leading to lower Peak Signal-to-Noise Ratio (PSNR) values in INRs. Following this insight, recent works (Tancik et al., 2020; Bai et al., 2023; Shi et al., 2024a; Cai et al., 2024) suggest that MLPs with balanced eigenvalues of ${\bm{K}}$ are less affected by spectral bias and achieve better PSNR values.
3.2 Improving spectral bias via Inductive Gradient Adjustment
@@
As previously discussed, eigenvalues of ${\bm{K}}$ almost govern convergence rates of MLPs to different components, and uneven spectrum of ${\bm{K}}$ leads to slow convergence of high frequency components. It is natural to adjust the spectrum of ${\bm{K}}$ purposefully to tailor the improvement of spectral bias. Following this intuition, we focus on the adjustment of eigenvalues of ${\bm{K}}$. This can be naturally achieved by constructing a transformation matrix ${\bm{S}}$ that has the same eigenvectors as ${\bm{K}}$ as follows:
$\Theta_{t+1}=\Theta_{t}-\eta\nabla_{\Theta}f({\bm{X}};\Theta){\bm{S}}{\bm{r}}_{t},$ (2)
where ${\bm{S}}=\sum_{i=1}^{N}(g_{i}(\lambda_{i})/\lambda_{i}){\bm{v}}_{i}{\bm{v}}_{i}^{\top}$; $\{g_{i}(\lambda_{i})\}_{i=1}^{N}$ denotes the desired spectrum. Geifman et al. (2023) theoretically analyze spectral impacts of equation 2 on ${\bm{K}}$ and validate it using a toy MLP with limited synthetic data.
@@
However, equation 2 is practically infeasible in a broad spectrum of INRs tasks. The first challenge is that an analytical expression for ${\bm{K}}$ is difficult to derive (Xu et al., 2021; Novak et al., 2022; Wang et al., 2023; Shi et al., 2024b). Specifically, the analytical expression for $\mathbb{E}_{f\sim\mathcal{N}(0,{\bm{\Sigma}})}[\dot{\sigma}(f(x))\dot{\sigma}(f(x^{\prime}))]$, which is the key of ${\bm{K}}$, becomes intractable as the depth of the network increases. The second challenge is that the size of the matrix ${\bm{K}}$ grows quadratically with the increase of the number of data points (Mohamadi et al., 2023). Considering Kodak image fitting task, it requires storing and decomposing a matrix containing about 1 trillion entries, which takes over eight terabytes of memory if stored in double precision.
To overcome the first challenge, we try to adjust eigenvalues of empirical NTK (eNTK) matrix $\tilde{{\bm{K}}}$ as $\tilde{{\bm{K}}}$ inherently exists in the linear dynamics model of MLPs with arbitrary width. Despite numerous studies (Xu et al., 2021; Novak et al., 2022; Mohamadi et al., 2023) exploring the connection between $\tilde{{\bm{K}}}$ and ${\bm{K}}$, impacts on spectral bias by $\tilde{{\bm{K}}}$-based adjustment are not evidently equivalent to ${\bm{K}}$-based adjustment. Through our theoretical analysis in Theorem 4.1, we prove that projection directions of $\tilde{{\bm{K}}}$ as well as eigenvalues converge to the corresponding parts of ${\bm{K}}$ as the network width increases. By this theoretical result, impacts on spectral bias by $\tilde{{\bm{K}}}$-based adjustment can be approximately equivalent to that of ${\bm{K}}$-based adjustment. We construct the transformation matrix $\tilde{{\bm{S}}}$ based on $\tilde{{\bm{K}}}$ and adjust the gradients as follows:
$\Theta_{t+1}$ $=\Theta_{t}-\eta\nabla_{\Theta_{t}}f({\bm{X}};\Theta_{t})\tilde{{\bm{S}}}{\bm{r}}_{t},$ (3)
where $\tilde{{\bm{S}}}=\sum_{i=1}^{N}(g_{i}(\tilde{\lambda_{i}})/\tilde{\lambda_{i}})\tilde{{\bm{v}}_{i}}\tilde{{\bm{v}}_{i}}^{\top}$; $\tilde{{\bm{K}}}=\nabla_{\Theta_{t}}f({\bm{X}};\Theta_{t})^{\top}\nabla_{\Theta_{t}}f({\bm{X}};\Theta_{t})=\sum_{i=1}^{N}\tilde{\lambda_{i}}\tilde{{\bm{v}}}_{i}\tilde{{\bm{v}}}_{i}^{\top}$.

### v2

@@ Abstract
Implicit Neural Representations (INRs), as a versatile representation paradigm, have achieved success in various computer vision tasks. Due to the spectral bias of the vanilla multi-layer perceptrons (MLPs), existing methods focus on designing MLPs with sophisticated architectures or repurposing training techniques for highly accurate INRs. In this paper, we delve into the linear dynamics model of MLPs and theoretically identify the empirical Neural Tangent Kernel (eNTK) matrix as a reliable link between spectral bias and training dynamics. Based on this insight, we propose a practical Inductive Gradient Adjustment (IGA) method, which could purposefully improve the spectral bias via inductive generalization of eNTK-based gradient transformation matrix. Theoretical and empirical analyses validate impacts of IGA on spectral bias. Further, we evaluate our method on different INRs tasks with various INR architectures and compare to existing training techniques. The superior and consistent improvements clearly validate the advantage of our IGA. Armed with our gradient adjustment method, better INRs with more enhanced texture details and sharpened edges can be learned from data by tailored impacts on spectral bias. The codes are available at: https://github.com/LabShuHangGU/IGA-INR.
Machine Learning, ICML
@@ 1 Introduction
The main idea of implicit neural representations (INRs) is using neural networks such as multi-layer perceptrons (MLPs) to parameterize discrete signals in an implicit and continuous manner. Benefiting from the continuity and implicit nature, INRs have gained great attention as a versatile signal representation paradigm (Luo et al., 2025) and achieved state-of-the-art performance across various computer vision tasks such as signal representation (Sitzmann et al., 2020; Liang et al., 2022; Saragadam et al., 2023; Kazerouni et al., 2024; Zhang et al., 2025), 3D shape reconstruction (Zhu et al., 2024; Zhang et al., 2024), and novel view synthesis (Mildenhall et al., 2021; Müller et al., 2022).
@@
However, obtaining high-precision INRs is non-trivial. MLPs with ReLU activations often fail to represent high-frequency details. Such tendency of MLPs to represent simple patterns of target function is referred to as spectral bias (Xu, 2018; Rahaman et al., 2019). To improve the performance of INRs, great efforts have been made to alleviate or overcome this bias of MLPs. These efforts have primarily focused on how MLPs are constructed, such as complex input embeddings (Tancik et al., 2020; Müller et al., 2022; Xie et al., 2023) and sophisticated activation functions (Sitzmann et al., 2020; Ramasinghe & Lucey, 2022; Saragadam et al., 2023; Zhu et al., 2024; Shenouda et al., 2024). Recently, two studies find that training dynamics adjustments based on traditional training techniques, i.e., Fourier reparameterized training (FR) (Shi et al., 2024a) and batch normalization (BN) (Cai et al., 2024), can overcome the spectral bias. Compared to the above approaches emphasizing architectural refinements, these methods could offer considerable representation performance gains while barely increasing the complexity of inference structures.
@@
Despite these works serendipitously revealing the great potential of training dynamics in alleviating spectral bias for better INRs, the underlying mechanism driving their improvements remains unclear. Moreover, there is no clear guidance on the choice of Fourier bases matrix for reparameterization or batch normalization layer to confront spectral bias with varying degrees in various INR tasks. Lack of reliable insight and guidance results in unstable and suboptimal improvements across different scenarios and INR models (Essakine et al., 2024), hindering their broader applications.
Notably, a stream of works (Ronen et al., 2019; Tancik et al., 2020; Shi et al., 2024a; Cai et al., 2024; Zhu et al., 2024) adopts the spectrum of the Neural Tangent Kernel (NTK) matrix (Jacot et al., 2018) to analyze spectral bias. Taking a step further, Geifman et al. (2023) have made a first attempt to modify the spectrum of the NTK matrix to improve convergence speeds, but limited to the toy MLP and restricted synthetic data. These prior works imply the potential of the NTK matrix to guide adjustments in training dynamics for better INRs. However, the NTK matrix is known for the general lack of an analytical expression and the quadratic growth in memory consumption as the data size increases (Xu et al., 2021; Novak et al., 2022; Mohamadi et al., 2023), especially in INRs, which involve various activation functions and dense predictions. Thus, how to effectively adjust training dynamics of MLPs to purposefully overcome spectral bias in INRs remains an open and valuable problem.
@@
In this paper, we delve into the linear dynamics model of MLPs and propose an effective training dynamics adjustment strategy guided by theoretical derivation to purposefully alleviate the spectral bias issue. Our strategy allows for tailored impacts on spectral bias, achieving more stable improvements. Specifically, these impacts on spectral bias are tailored by purposefully adjusting the spectrum of the NTK matrix, which closely connects the spectral bias with training dynamics. Given that the NTK matrix is not available in most cases, we theoretically justify the empirical NTK (eNTK) matrix (Novak et al., 2022) as a tractable surrogate. Empirical evidence corroborates these theoretical results. Based on this insight, we further propose an inductive gradient adjustment (IGA) method, which could purposefully improve the spectral bias via inductive generalization of the eNTK-based gradient transformation matrix with millions of data points. We give both theoretical and empirical analysis of our method, demonstrating how it tailors impacts on spectral bias. Besides, our IGA method can work with previous structural improvement methods such as Positional Encoding (Tancik et al., 2020) and periodic activations (Sitzmann et al., 2020). We validate our method in various vision applications of INRs and compare it to previous training dynamics adjustment methods. Experimental results show the superiority of our IGA method. Our contributions are summarized as follows:
@@
  • We connect spectral bias with linear dynamics model of MLPs and derive a training dynamics adjustment strategy guided by a theoretical standpoint to purposefully improve the spectral bias.
@@
  • We present comprehensive experimental analyses on synthetic data and a range of implicit neural representation tasks. Compared to existing training dynamics approaches, our method enables tailored improvements in mitigating the spectral bias of standard MLPs, resulting in implicit neural representations with enhanced high-frequency details.
@@ 2 Related Work
Implicit neural representations. Implicit neural representations (INRs) representing the discrete signal as an implicit continuous function by MLPs have gained lots of attention. Comparing to traditional grid-based methods, INRs offer remarkable accuracy and memory efficiency across various representation tasks such as 1D audio representation (Kim et al., 2022), 2D image representation (Klocek et al., 2019; Strümpler et al., 2022), 3D shape representation (Park et al., 2019; Martel et al., 2021), novel view synthesis (Mildenhall et al., 2021; Saratchandran et al., 2024). However, vanilla MLPs with ReLU activations fail to represent complex signals. Therefore, various modifications have been studied. One category focuses on input embedding, such as using Positional Encoding (PE) (Mildenhall et al., 2021) or learned features (Takikawa et al., 2021; Martel et al., 2021; Xie et al., 2023). A different category of modifications concentrates on better activations; for instance, periodic activations (Sitzmann et al., 2020; Fathony et al., 2020; Zhu et al., 2024) and Gabor wavelet activations (Saragadam et al., 2023) provide more accurate and robust representations. Considering a broader framework, these works can be summarized as efforts to enhance the representational capacity of INRs or MLPs (Yüce et al., 2022). Recently, another category, distinct from those previously discussed, focuses on the training process of MLPs. Shi et al. (2024a) find that learning parameters in the Fourier domain, i.e., Fourier reparameterized training (FR), can improve approximation accuracy of INRs without altering the inference structure. Cai et al. (2024) propose that the classic batch normalization layer (BN) can also improve the performance of INRs.
@@
Spectral bias. Spectral bias refers to the learning bias of MLPs to prioritize learning simple patterns or low-frequency components of the target function. Great efforts have been made to dissect this bias. Rahaman et al. (2019) find that spectral bias varies with the width, depth and input manifold of MLPs. Xu (2018) attribute this bias of MLPs with Tanh activation to the uneven distribution of gradients in the frequency domain. From the perspective of linear dynamics model of MLPs, spectral bias is induced by the uneven spectrum of the corresponding Neural Tangent Kernel (NTK) matrix (Jacot et al., 2018; Arora et al., 2019; Ronen et al., 2019). Inspired by this theoretical result, a series of works (Tancik et al., 2020; Shi et al., 2024a; Cai et al., 2024) try to elucidate spectral bias by observing spectral distribution of the NTK matrix and find that PE, FR and BN help to attain more even spectrum. In this paper, we adopt the insight of attaining more even spectrum of the NTK matrix. For the first time, we show that modifying the eNTK matrix spectrum can have the similar impacts on spectral bias and propose a practical gradient adjustment method, which improves spectral bias of INRs via inductive generalization of eNTK-based gradient transformation matrix.
@@ 3 Method
In this section, we first review the linear dynamics model of MLPs, indicating the potential of adjusting the NTK spectrum to overcome spectral bias in INRs. Then we analyze the NTK-based adjustment and show that the NTK matrix’s intractability renders it impractical. By theoretical analysis, we justify the eNTK matrix as a tractable proxy, but it succumbs to dimensionality challenges as data size grows. Thus, we propose a practical gradient adjustment method that improves spectral bias via inductive generalization of eNTK-based gradient transformation matrix.
Figure 1: Evolution of approximation error with training iterations in time and Fourier domain. Line plots visualize MSE loss curves of MLPs with $1024$ and $4096$ neurons optimized by four gradient adjustments, i.e., ${\bm{I}},{\bm{S}},\tilde{{\bm{S}}},\tilde{{\bm{S}}}_{e}$. The shaded area of lines indicates training fluctuations. Heatmaps show the relative error $\Delta_{k}$ in Eq. 6 on four frequency bands. Details are in the experiment 1 of Sec. 4.
3.1 Connecting spectral bias with training dynamics
Without loss of generality and for simplicity, we consider a scalar discrete signal. We define the vector of signal values as ${\bm{y}}=(y_{1},\dots,y_{N})\in\mathbb{R}^{N}$, and the corresponding coordinate vectors $({\bm{x}}_{1},\dots,{\bm{x}}_{N})^{\top}$, denoted by ${\bm{X}}\in\mathbb{R}^{N\times d}$. The INR of this signal is parameterized as $f({\bm{X}};\Theta)$, where $f(\cdot;\Theta)$ denotes an MLP. Parameters $\Theta\in\mathbb{R}^{q}$ are optimized with respect to $\frac{1}{2}\sum_{i=1}^{N}(f({\bm{x}}_{i};\Theta)-y_{i})^{2}$ via gradient descent. The residual $\left(f({\bm{x}}_{i};\Theta_{t})-y_{i}\right)_{i=1}^{N}$ at time step $t$ is denoted as ${\bm{r}}_{t}$. Training dynamics of ${\bm{r}}_{t}$ can be approximated by the linear dynamics model: ${\bm{r}}_{t}=({\bm{I}}-\eta{\bm{K}}){\bm{r}}_{t-1}$ under the conditions of wide networks and small learning rates $\eta$ (Du et al., 2018; Arora et al., 2019; Lee et al., 2019). ${\bm{K}}$ is the NTK matrix defined by $\mathbb{E}_{\Theta_{0}}[\nabla_{\Theta}f({\bm{X}};\Theta_{0})^{\top}\nabla_{\Theta}f({\bm{X}};\Theta_{0})]$, where $\nabla_{\Theta}f({\bm{X}};\Theta_{0})\in\mathbb{R}^{q\times N}$. With the eigenvalue decomposition of ${\bm{K}}=\sum_{i=1}^{N}\lambda_{i}{\bm{v}}_{i}{\bm{v}}_{i}^{\top}$ and recurrence unfolding of the linear dynamics model, training dynamics of ${\bm{r}}_{t}$ are characterized as follows (Arora et al., 2019):
@@ 3.1 Connecting spectral bias with training dynamics
Eq. 1 shows that convergence rate of $f({\bm{x}},\Theta)$ at the projection direction ${\bm{v}}_{i}^{\top}$ with larger $\lambda_{i}$ is faster. For vanilla MLPs, projection directions related to high frequencies are consistently assigned to small eigenvalues, while those related to low frequencies correspond to larger eigenvalues (Ronen et al., 2019; Bietti & Mairal, 2019; Heckel & Soltanolkotabi, 2020). Such an uneven spectrum of ${\bm{K}}$ leads to extremely slow convergence rates to high frequency components of the signal, leading to a reduction in average approximation accuracy (e.g., MSE loss) as well as frequency-aware metrics (e.g., SSIM, MS-SSIM (Wang et al., 2003)) in INRs. Following this insight, recent works (Tancik et al., 2020; Bai et al., 2023; Shi et al., 2024a; Cai et al., 2024) suggest that MLPs with a more even spectrum of ${\bm{K}}$ are less affected by spectral bias.
3.2 Inductive Gradient Adjustment
@@
As previously discussed, such an uneven spectrum of ${\bm{K}}$ leads to slow convergence of high frequency components. Naturally, adjusting the evenness of the spectrum of ${\bm{K}}$ is expected to tailor the improvement of spectral bias. Following this intuition, we focus on the adjustment of eigenvalues of ${\bm{K}}$, i.e., ${\bm{K}}$-based adjustment. Inspired by the theoretical insights of Geifman et al. (2023), this can be achieved by multiplying a transformation matrix ${\bm{S}}$ as follows [^1] This matrix is also known as the pre-conditioned matrix in the field of optimization.:
$\vspace{-0.5em}\Theta_{t+1}=\Theta_{t}-\eta\nabla_{\Theta}f({\bm{X}};\Theta){\bm{S}}{\bm{r}}_{t},$ (2)
where ${\bm{S}}=\sum_{i=1}^{N}(g_{i}(\lambda_{i})/\lambda_{i}){\bm{v}}_{i}{\bm{v}}_{i}^{\top}$; $\{g_{i}\}_{i=1}^{N}$ denote the potential eigenvalue transformations. As driven by the training dynamics of Eq. 2, the spectrum of the modified kernel matrix ${\bm{K}}$ can be theoretically approximated as $\{g_{i}(\lambda_{i})\}_{i=1}^{N}$.
@@
Although Eq. 2 offers a theoretical approach to adjust impacts of training dynamics on spectral bias, Eq. 2 is almost infeasible in practical INRs tasks. The first challenge is that an analytical expression for ${\bm{K}}$ is difficult to derive (Xu et al., 2021; Novak et al., 2022; Wang et al., 2023; Shi et al., 2024b). Specifically, the analytical expression $\mathbb{E}_{\Theta_{0}}[\nabla_{\Theta}f({\bm{X}};\Theta_{0})^{\top}\nabla_{\Theta}f({\bm{X}};\Theta_{0})]$ of ${\bm{K}}$ becomes intractable as the depth of the network increases. The second challenge is that the size of ${\bm{K}}$ grows quadratically with the number of data points (Mohamadi et al., 2023). Considering the Kodak image fitting task (Saragadam et al., 2023; Shi et al., 2024a), it requires decomposing a matrix on the order of $10^{11}$ entries, which occupies over $8192$GiB (gibibytes) of memory if stored in double precision.
To overcome the first challenge, numerous studies (Xu et al., 2021; Novak et al., 2022; Mohamadi et al., 2023) have studied the eNTK matrix $\tilde{{\bm{K}}}$ defined by $\nabla_{\Theta_{t}}f({\bm{X}};\Theta_{t})^{\top}\nabla_{\Theta_{t}}f({\bm{X}};\Theta_{t})$ and Geifman et al. (2023) show that replacing ${\bm{K}}$ with $\tilde{{\bm{K}}}$ ($\tilde{{\bm{K}}}$-based) in Eq. 2 can accelerate convergence on toy settings. However, whether $\tilde{{\bm{K}}}$-based adjustment has the similar impacts on spectral bias as in Eq. 2 and remains consistently effective in general settings, is an open question. To justify feasibility, we prove that eigenvectors of $\tilde{{\bm{K}}}$ as well as eigenvalues converge to the corresponding parts of ${\bm{K}}$ as the network width increases. Namely, impacts on spectral bias by $\tilde{{\bm{K}}}$-based adjustment can be approximately equivalent to that of ${\bm{K}}$-based adjustment. The detailed theoretical analysis is presented in our Theorem 3.1. Then, we empirically validate the aforementioned equivalence and demonstrate that $\tilde{{\bm{K}}}$-based adjustment can tailor the improvement of spectral bias on more general settings in Sec. 4. The formula of $\tilde{{\bm{K}}}$-based adjustment is as follows:
$\vspace{-0.5em}\Theta_{t+1}=\Theta_{t}-\eta\nabla_{\Theta_{t}}f({\bm{X}};\Theta_{t})\tilde{{\bm{S}}}{\bm{r}}_{t},$ (3)
where $\tilde{{\bm{K}}}=\sum_{i=1}^{N}\tilde{\lambda}_{i}\tilde{{\bm{v}}}_{i}\tilde{{\bm{v}}}_{i}^{\top}$; $\tilde{{\bm{S}}}=\sum_{i=1}^{N}(g_{i}(\tilde{\lambda}_{i})/\tilde{\lambda}_{i})\tilde{{\bm{v}}}_{i}\tilde{{\bm{v}}}_{i}^{\top}$.